OpenStudy (anonymous):

Fan + Medal Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (astrophysics):

\[x = rcos \theta\] \[y = rsin \theta\] For polar to cartesian coordinates. For cartesian to polar: \[r^2 = x^2+y^2 ~~~~ \theta = \tan ^{-1} \left( \frac{ y }{ x } \right)\]

OpenStudy (astrophysics):

Note you want r so \[r = \sqrt{x^2+y^2}\]

OpenStudy (anonymous):

r = sqrt(8) or 2.828

OpenStudy (astrophysics):

Sure, now find theta

OpenStudy (anonymous):

theta = 45

OpenStudy (anonymous):

now how do we find which quadrant it is in?

OpenStudy (astrophysics):

Yeah but notice  theta gives us the value for the first and we need the fourth quadrant.

OpenStudy (astrophysics):

Nice, you figured that out

OpenStudy (anonymous):

oh that makes sense

OpenStudy (anonymous):

so 315

OpenStudy (astrophysics):

Yeah that seems reasonable so 7pi/4

OpenStudy (anonymous):

which is \[\large(\frac{ \sqrt{2} }{ 2 }, \frac{ \sqrt{-2} }{ 2 })\]

OpenStudy (anonymous):

on the unit circel

OpenStudy (anonymous):

circle*

OpenStudy (astrophysics):

When you're looking for theta, you need unit circle :)

OpenStudy (astrophysics):

Seems you got a good grasp on it though, nice work.

OpenStudy (astrophysics):

\[\sqrt{8} \implies 2\sqrt{2}\]

OpenStudy (anonymous):

so what is the answer exactly? I need 2 points in Polar Coordinates of the original thing

OpenStudy (anonymous):

these are my options: (2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 135°), (-2 square root of 2, 315°) (2 square root of 2, 315°), (-2 square root of 2, 135°) (2 square root of 2, 45°), (-2 square root of 2, 225°)

OpenStudy (astrophysics):

Oh it wants two pairs

OpenStudy (astrophysics):

So now what do you think the other will be?

OpenStudy (anonymous):

Im blanking @Astrophysics

OpenStudy (anonymous):

I think it's the third option. Am I right @zepdrix ?

OpenStudy (astrophysics):

Oh I see a mistake

OpenStudy (astrophysics):

What does \[\theta = \tan^{-1}\left( \frac{ -2 }{ 2 } \right)\] give?

zepdrix (zepdrix):

So you found your first polar pair by rotating 315, and then extending outward 2sqrt2? Looks good.